OpenStudy (anonymous):

What is the range of the graph of y = (x – 4)^2 + 3? y < 4 y >4 y >3 y <3 They have lines under them

OpenStudy (anonymous):

I think it's 4 but I don't know which one HELP anyone Please

OpenStudy (anonymous):

I don't understand this question!!

jimthompson5910 (jim_thompson5910):

Hint: In general, the range of \[\Large y = a(x-h)^2 + k\] is \[\Large y \ge k\]

OpenStudy (anonymous):

first of all "range" is a range of numbers, not one number

jimthompson5910 (jim_thompson5910):

if 'a' is greater than 0

OpenStudy (anonymous):

so it's y>4

jimthompson5910 (jim_thompson5910):

What is the value of k in this case?

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

no, that's the value of h

jimthompson5910 (jim_thompson5910):

Use \[\Large y = a(x-h)^2 + k\] as your template

OpenStudy (anonymous):

3

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

since a > 0, we know that the range is then \[\Large y \ge 3\]

jimthompson5910 (jim_thompson5910):

since in general the range is \[\Large y \ge k\] when a > 0

OpenStudy (anonymous):

Really I thought it was 4

jimthompson5910 (jim_thompson5910):

nope, definitely not 4

OpenStudy (anonymous):

okay thannks for your help

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

i have a question do you know what "range" means? if so, then this is straightforward if not it is probably very confusing

jimthompson5910 (jim_thompson5910):

that's where the min value occurs

jimthompson5910 (jim_thompson5910):

that doesn't say anything about the range

jimthompson5910 (jim_thompson5910):

if you do this instead http://www.wolframalpha.com/input/?i=range+y+%3D+%28x+%E2%80%93+4%29^2+%2B+3 you'll get what you're looking for But it's always good to know how to do this without a calculator

OpenStudy (anonymous):

ohh okay i feell stupid!!

jimthompson5910 (jim_thompson5910):

happens all the time, so no worries